• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 17:15:48 on November 10, 1998

    Entry number 5119

    keyword=Beam Time Accounting
    Hall A

    11/10/1998 0800-1600
    Experiment #: E94-010
    Shift Leader: B. Hersman
    Shift Workers: C. Jutier, R. Suleiman, A. Saha, S. Incerti

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    0 0 0 0 0 0 0.75 1.0 1.75
    Beam Available
    but Not in Use
    0 0 0 0 0 0.75 0.25 0 1
    Beam Not Available
    or Unacceptable
    1.0 1.0 1.0 1.0 1.0 0.25 0 0 5.25
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1.0 1.0 1.0 1.0 1.0 0.25 0.75 1.0 7
    Planned Config-
    uration Change
    0 0 0 0 0 0.75 0.25 0 1
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: - MCC cleaned the orbit, did B-Scope and we did harp scan and raster check. - NMR at 13:50 gives 32.6 % (H2O)

    Data Entry by: B. Hersman